# A Power Method for Computing Singular Value Decomposition

### v2

@@ 1.1 Classical SVD computation
The Frobenius norm of the matrix $\mathbf{X}$ is:
@@
where $\operatorname{tr}[.]$ is the matrix trace operator. Let $\mathbf{X}_{r}$ be the matrix obtained using the first $r$ terms in the sum of relation (2). The Frobenius norm allows to quantify the error of the $r$ rank approximation for $\mathbf{X}$, [15, 18]:
@@
The SVD also allows to solve the linear system of equations which occurs in many engineering problems: $\mathbf{Ax}=\mathbf{b}$, where $\mathbf{A}$ and $\mathbf{b}$ are known matrix and vector of size $m\times n$ and $n$, respectively. The entries of the unknown vector $\mathbf{x}$ should be found. For a square matrix, $\mathbf{A}$, the Lanczos algorithm, [23], allows to obtain the solution through matrix eigenvalues, see below. For a non-square matrix $\mathbf{A}$, the normal equation, $\mathbf{A}^{\mathsf{T}}\mathbf{Ax}=\mathbf{A}^{\mathsf{T}}\mathbf{b}$ which involves a symmetric can be used to search for the vector $\mathbf{x}$. For a sparse matrix with large $m$ and $n$ values, the Lanczos algorithm has been extended by Larsen, [25], to bidiagonalization of the non-square matrix leading to perform its SVD. For high dimensional data, when the purpose is to get an approximation of the data matrix, it may not be necessary to get all singular values. The Lanczos algorithm extension proposed in [1] allows to compute the few first largest or smallest singular values and their associated vectors.
1.2 Probabilistic SVD computation
Performing complete SVD can be computationally costly, especially for high dimensional data. When the first important singular values are required, the randomized methods may be used. The computation of the SVD for massive data has been considered in [27, 17, 16, 38], where, only the needed $r$ singular values and their associated vectors are computed using a probabilistic (randomized) approach. The search consists of finding column orthonormal matrix $\mathbf{Q}$ of size $m\times r$ such that:
$\|\mathbf{X}-\mathbf{QQ}^{\mathsf{T}}\mathbf{X}\|\leq\epsilon$ (6)
[…]
This relation shows a ($12q+$ 1)-power of the singular values for this matrix $\mathbf{Y}$. The rapid singular values decay leads to a fast convergence algorithm, [17, 16]. The paper [17] gives historical overview of the randomized computing and its applications in linear algebra problems. Using the probabilistic SVD method, the rank $r$ approximation error is slightly greater than the bound given in relation (5).
1.3 Eigendecomposition
The eigendecomposition can be performed for any square matrix. Here, only the symmetric matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ case is considered. The eigendecomposition of this matrix is:
where $\boldsymbol{\Lambda}$ is a diagonal matrix which entries are nonnegative-valued scalars or eigenvalues: $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}\geq 0$, $\mathbf{V}$ is an orthogonal matrix, $\mathbf{V}^{\mathsf{T}}\mathbf{V}=\mathbf{I}_{n}$, which columns are the eigenvectors associated with the eigenvalues. Like in relation (2), the eigendecomposition can be expressed as the sum of rank one matrices or spectral decomposition:
where $\mathbf{v}_{j}$ is the $j$-th column vector of the matrix $\mathbf{V}$. The above sum can also be truncated to get an lower rank $r<n$ approximation for the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. The column orthogonal property of the matrix $\mathbf{V}$ allows to have:
where $q$ is an integer, $q\geq 1$. That means, all power $q$ of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors, only the eigenvalues differ by the power $q$ and we also have: $\lambda_{1}^{q}\geq\lambda_{2}^{q}\geq\ldots\geq\lambda_{n}^{q}\geq 0$. At least, the first eigenvalue is nonzero. Therefore, we have:
$(a)$ ::: $1\geq(\lambda_{2}/\lambda_{1})\geq(\lambda_{3}/\lambda_{1})\geq\ldots\geq(\lambda_{n}/\lambda_{1})$ (11)
$(b)$ ::: $1\geq(\lambda_{2}/\lambda_{1})^{q}\geq(\lambda_{3}/\lambda_{1})^{q}\geq\ldots\geq(\lambda_{n}/\lambda_{1})^{q}$ (12)
$(c)$ ::: $(\lambda_{j}/\lambda_{1})^{q}\leq(\lambda_{j}/\lambda_{1})\mbox{ ; }j=1,2,\ldots,n$ (13)
In the principal component analysis, see below, the columns of the matrix $\mathbf{V}$ correspond to projection subspaces. One can observe that $-\mathbf{V}$ is also solution of the eigendecomposition in relation (8). This observation allows to fulfil the Perron-Frobenius theorem, [31, 12], which stipulates that the entries of the main eigenvector of a nonnegative-valued matrix are all positive or zero. Nonnegative-valued matrices occur for many real-life data, e.g. genes expression and images. The main eigenvector is used in ranking applications, [5, 24]. To compute the eigenvalue with maximum modulus and the associated eigenvector, the iterative power method, [15], is very often used, see algorithm 2 in the Annexes.
The convergence rate of the power method depends on the modulus of the second eigenvalue compared to the first, [15, 7]. At the convergence, the eigenvector is used to get the associated eigenvalue: $\mathbf{X}^{\mathsf{T}}\mathbf{Xv}_{j}=\lambda_{j}\mathbf{v}_{j}$. Thank to relation (9), the power method can be used to obtain all eigenvectors and their associated eigenvalues. Indeed, let us define $\mathbf{S}_{1}=\mathbf{X}^{\mathsf{T}}\mathbf{X}$. When the first eigenvector is obtained, a new matrix can be formed as $\mathbf{S}_{2}=\mathbf{S}_{1}-\lambda_{1}\mathbf{v}_{1}\mathbf{v}_{1}^{\mathsf{T}}$. The matrix $\mathbf{S}_{2}$ is used to get the second eigenvector and associated eigenvalue. This process is repeated to get all the eigenvector and eigenvalue pairs. That is the Jordan SVD algorithm, [37]. For a rank deficient matrix $\mathbf{X}$, the number of eigenvalues is less than $n$. The power method proposed below allows to compute all required eigenvalues together.
2 Computing singular values using a power method
Given a matrix $\mathbf{X}$ of size $m\times n$, $m\geq n$, we first show the link between it’s singular value decomposition factors and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. Using the expression (1) of $\mathbf{X}$ in the relation (8) allows to write:
Comparing the above expression to relation (8) allows to express the SVD factors using the eigendecomposition factors:
[…]
@@ 2 Computing singular values using a power method
The entries of the diagonal matrix $\boldsymbol{\Lambda}$ are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$. Relation (14) shows a one-to-one relationship between the SVD factors of the matrix $\mathbf{X}$ and the eigendecomposition factors of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. From the above relation, the SVD of the matrix $\mathbf{X}$ can be written as:
@@
When $\mathbf{W}\triangleq\mathbf{Q}$, the above relation is equivalent to that is used in relation (6) and is obtained when performing the eigendecomposition of the matrix $\mathbf{XX}^{\mathsf{T}}$. Since we assumed that $n\leq m$, working with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ necessitates less computational load compared to the matrix $\mathbf{XX}^{\mathsf{T}}$. To obtain the column orthogonal matrix $\mathbf{W}$ given $\mathbf{X}$, a norm of the difference between the left-hand and right-hand side terms in the above relation can be minimized. In the probabilistic SVD approach, a column orthonormal matrix of size $r+\ell$, $\ell>r$, is used and the QR decomposition is performed followed by matrices multiplication to get the SVD result. Here, the entries of a random column orthogonal matrix of size $n\times r$ are iteratively modified to fit the factor $\mathbf{W}$ for the SVD result of the matrix $\mathbf{X}$. The algorithm used is similar to the power method but all eigenvectors are simultaneously recovered at a time. A new parameter is introduced to increase the convergence rate, compared to the basic power method.
2.1 Search for the projection matrix using the gradient method
To obtain the matrix $\mathbf{W}$ in relation (15), the Frobenius norm of the difference between $\mathbf{X}$ and $\mathbf{XWW}^{\mathsf{T}}$ is minimized, subject to the column orthogonality constraint for the matrix $\mathbf{W}$. The following constrained optimization problem is solved:
$\min\mathsf{J}(\mathbf{W})\mbox{, }\mathsf{J}(\mathbf{W})=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}\mbox{ subject to: }\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}$ (16)
where .$\|.\|_{\mathsf{F}}^{2}$ is the square of the Frobenius norm. The Lagrangian of this optimization problem is:
@@ 2.1 Search for the projection matrix using the gradient method
where the entries of the matrix $\Upsilon$ are the Lagrange multipliers associated with the equality constraints. To minimize $\mathsf{J}(\mathbf{W})$, the gradient descent method can be used with the column orthogonal constraint for the matrix $\mathbf{W}$. For the gradient search optimization method, the solution at iteration $t$ is obtained after updating the previous one: $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla\mathcal{L}(\mathbf{W},\Upsilon)$, where $\eta$ is a positive step parameter and $\nabla\mathcal{L}(\mathbf{W},\Upsilon)$ is the gradient of the Lagrangian. To obtain the matrix $\mathbf{W}$ leading to minimum of the Lagrangian, the derivative of $\mathcal{L}(\mathbf{W},\Upsilon)$ with respect to $\mathbf{W}$ and $\Upsilon$ should be computed.
@@
For any column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$, we then have:
@@
$\left\|\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right\|_{\mathsf{F}}^{2}$ $=$ 00 0 (19)
For an $n$-order orthogonal matrix $\mathbf{W}$, we also have $\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{I}_{n}$. However, all $n$-order orthogonal matrices do not allow to verify relation (8) for a given matrix $\mathbf{X}$. In addition, we would like to be able to compute a column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $r<n$. In this case, we still have $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$ but $\mathbf{W}\mathbf{W}^{\mathsf{T}}\neq\mathbf{I}_{n}$. The matrix $\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}$ is symmetric and has nonnegative-valued eigenvalues. The trace of this matrix is obtained as the sum of its eigenvalues. For an $n$-order column orthogonal matrix $\mathbf{W}$, relation (18) shows that the global minimum of the objective function $\mathsf{J}(\mathbf{W})$ is zero. For an $r$-column orthogonal matrix $\mathbf{W}$, $r<n$, we have:
@@
where $\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}$ is the diagonal matrix which entries are the $r$ first singular values, see the text under relation (14).
The relation (118) in Petersen & Petersen [32, page 13], can be written as:
@@
The matrix $\mathbf{W}$ in the above relation is that available in the previous iteration: $\mathbf{W}^{(t-1)}$. Then, the gradient updating equation is: $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{W}^{(t-1)}$. Therefore, at iteration $t=1,2,\ldots$ and given column orthogonal matrix $\mathbf{W}^{(t-1)}$, the gradient descent optimization search method consists of two alternate steps to obtain the update:
@@
$\mathbf{W}^{(t)}$ $=$ $Gram-Schmidt(\tilde{\mathbf{W}})$ (21)
where $\mathbf{G}\triangleq(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$, $\eta$ is a positive step parameter and $q\geq 1$ is an integer. The settings for $\eta$ and $q$ are discussed later. The method to use for obtaining the column orthogonal matrix $\mathbf{W}$ is summarized in algorithm 1.
Data: matrix $\mathbf{X}$ of size $m\times n$, ($m\geq n$), settings $\eta$, $q$ and $\epsilon$;
Result: matrix $\mathbf{W}$ with orthogonal columns;
1 Initialization: select randomly entries of a matrix $\mathbf{W}$ from a normal or an uniform distribution and perform orthogonalization $\mathbf{W}^{(0)}=$ Gram-Schmidt($\mathbf{W}$), compute $\mathbf{G}=(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$;
2 for $t\leftarrow 1,2,\ldots$ do
@@
Algorithm 1 Gradient search for columns orthonormal matrix $\mathbf{W}$
This algorithm is similar to the power method, see below.
2.1.1 Factors for the eigendecomposition of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and for the SVD of $\mathbf{X}$
The algorithm 1 leads to a column orthogonal matrix $\mathbf{W}$ that fits the matrix $\mathbf{X}$. Using relation (8), the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are the entries of the diagonal matrix $\boldsymbol{\Lambda}$. These eigenvalues are the diagonal entries of the matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$ which is symmetric. The diagonal matrix of the SVD is: $\mathbf{D}=\boldsymbol{\Lambda}^{1/2}$. Finally the matrix $\mathbf{U}$ of the SVD is: $\mathbf{U}=\mathbf{XW}\boldsymbol{\Lambda}^{-1/2}$.
2.1.2 Relation with the power method
Algorithm 1 is in fact the power method allowing to obtain all eigenvalues together and not one by one as done using the classical power method algorithm 2, see Annexes. The matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ is used in this algorithm instead of the matrix $\mathbf{A}\triangleq\mathbf{X}^{\mathsf{T}}\mathbf{X}$. These two matrices have the same eigenvectors. Indeed, the eigenvector $\mathbf{z}$ associated with the eigenvalue $\gamma$ of the matrix $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ allows to write:
$(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{z}=\gamma\mathbf{z}\Rightarrow\mathbf{X}^{\mathsf{T}}\mathbf{Xz}=\left(\frac{\gamma-1}{\eta}\right)\mathbf{z}$
The above relation means the matrices $\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)$ and $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors which are associated with related eigenvalues, $j=1,2,\ldots,n$:

### v4

@@ 2 Computing the SVD by using a power method
The proposed power method is based on the Frobenius norm of the matrix $\mathbf{X}$:
@@
where $\operatorname{tr}[.]$ is the matrix trace operator. Let $\mathbf{X}_{r}$ be the matrix obtained using the $r$ first terms in the sum of relation (2). The Frobenius norm allows to quantify the error of the $r$-rank $\mathbf{X}$ approximation, [24, 29]:
@@
Relation (6) shows a one-to-one relationship between the SVD factors of $\mathbf{X}$ and the EVD factors of $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. From relation (6), the SVD of the matrix $\mathbf{X}$ can be written as:
@@
When $\mathbf{W}\triangleq\mathbf{Q}$, the above relation is equivalent to that is used in relation (7) and is obtained when performing the EVD of the matrix $\mathbf{XX}^{\mathsf{T}}$. Since we assumed that $n\leq m$, working with the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ necessitates less computational load compared to the matrix $\mathbf{XX}^{\mathsf{T}}$. To obtain the column orthogonal matrix $\mathbf{W}$ given $\mathbf{X}$, a norm of the difference between the left-hand and right-hand side terms in the above relation can be minimized. In the probabilistic SVD approach, a column orthogonal matrix of size $\ell, >r\ell>r$, is used and the QR decomposition is performed followed by matrices multiplication to get the SVD. Here, the entries of a random column orthogonal matrix of size $n\times r$ are iteratively modified to fit the factor $\mathbf{W}$ of the SVD. The algorithm used is similar to the power method but the $r\leq n$ eigenvectors are simultaneously recovered.
2.1 Search for the projection matrix by using the gradient method
To obtain the matrix $\mathbf{W}$ in relation (15), the Frobenius norm of the difference between $\mathbf{X}$ and $\mathbf{XWW}^{\mathsf{T}}$ is minimized, subject to the column orthogonality constraint for the matrix $\mathbf{W}$.
$\min\mathsf{J}(\mathbf{W})\mbox{; }\mathsf{J}(\mathbf{W})=\frac{1}{2}\left\|\mathbf{X}-\mathbf{XWW}^{\mathsf{T}}\right\|_{\mathsf{F}}^{2}\mbox{ subject to: }\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}$ (16)
The Lagrangian of this optimization problem is:
@@ 2.1 Search for the projection matrix by using the gradient method
where the entries of the matrix $\Upsilon$ are the Lagrange multipliers associated with the equality constraints. To minimize $\mathsf{J}(\mathbf{W})$, the gradient descent method is used, the solution at iteration $t$ is obtained after updating the previous one: $\mathbf{W}^{(t)}=\mathbf{W}^{(t-1)}-\eta\nabla\mathcal{L}(\mathbf{W},\Upsilon)$, where $\eta$ is a positive step parameter and $\nabla\mathcal{L}(\mathbf{W},\Upsilon)$ is the gradient of the Lagrangian. The derivative of $\mathcal{L}(\mathbf{W},\Upsilon)$ with respect to $\mathbf{W}$ and $\Upsilon$ are then required.
@@
For any column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$, that allows to write:
@@
$\left\|\mathbf{W}^{\mathsf{T}}\mathbf{W}-\mathbf{I}\right\|_{\mathsf{F}}^{2}$ $=$ $0$ (19)
For any $n$-order orthogonal matrix $\mathbf{W}$, we have $\mathbf{W}\mathbf{W}^{\mathsf{T}}=\mathbf{I}_{n}$. However, all $n$-order orthogonal matrices do not allow to verify relation (4) for a given matrix $\mathbf{X}$. In addition, we would like to be able to compute a column orthogonal matrix $\mathbf{W}$ of size $n\times r$, $r<n$. In this case, we still have $\mathbf{W}^{\mathsf{T}}\mathbf{W}=\mathbf{I}_{r}$ but $\mathbf{W}\mathbf{W}^{\mathsf{T}}\neq\mathbf{I}_{n}$. The matrix $\mathbf{XX}^{\mathsf{T}}-\mathbf{XWW}^{\mathsf{T}}\mathbf{X}^{\mathsf{T}}$ is symmetric and has nonnegative-valued eigenvalues. The trace of this matrix is the sum of its eigenvalues. For a $n$-order column orthogonal matrix $\mathbf{W}$, relation (18) shows that the global minimum of the objective function $\mathsf{J}(\mathbf{W})$ is zero. For a $r$-column orthogonal matrix $\mathbf{W}$, $r<n$, we have:
@@
where $\mathbf{XW}(\mathbf{XW})^{\mathsf{T}}$ is the diagonal matrix which entries are the $r$ first singular values, see the text under relation (6). The relation (118) in [51, page 13], can be written as:
@@
The matrix $\mathbf{W}$ in the above relation is that available in the previous iteration, $\mathbf{W}^{(t-1)}$. Then, the gradient updating equation is: $(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X})\mathbf{W}^{(t-1)}$. Therefore, at iteration $t=1,2,\ldots$ and given column orthogonal matrix $\mathbf{W}^{(t-1)}$, two alternate steps are used to get the update $\mathbf{W}^{(t)}$:
@@
[…]
where $\mathbf{G}\triangleq\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$, $\eta$ is a positive step parameter and $q\geq 1$ is an integer. The settings for $\eta$ and $q$ are discussed later. The method to use for obtaining the column orthogonal matrix $\mathbf{W}$ is summarized in algorithm 1.
Data: matrix $\mathbf{X}$ of size $m\times n$, settings $\eta$, $q$, $\epsilon$ and maximum number of iterations (itmaxitmax);
Result: column orthogonal matrix $\mathbf{W}$;
1 Initialization: select randomly entries of a matrix $\mathbf{W}$ from a normal or an uniform distribution and perform orthogonalization $\mathbf{W}^{(0)}=$ Gram-Schmidt($\mathbf{W}$), compute $\mathbf{G}=\mathbf{I}_{n}+\eta(\mathbf{X}^{\mathsf{T}}\mathbf{X})^{q}$;
2 for $t\leftarrow 1,2,\ldots,itmax$ do
@@
Algorithm 1 Gradient search for column orthogonal matrix $\mathbf{W}$
The columns of the result matrix $\mathbf{W}$ are the eigenvectors matrix of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and the right singular vectors matrix of $\mathbf{X}$. Using relation (4), the eigenvalues of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ are the entries of the diagonal matrix $\boldsymbol{\Lambda}$ or the diagonal entries of the symmetric matrix $(\mathbf{XW})^{\mathsf{T}}\mathbf{XW}$. Relation (6) allows to get the SVD factor matrices $\mathbf{D}$ and $\mathbf{U}$.
2.2 Relation with the power method
Algorithm 1 is a power method allowing to get the $r\leq n$ eigenvectors simultaneously and not one by one like by the classical power method in algorithm 2, see Annexes. The matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ is used in algorithm 1 instead of the matrix $\mathbf{X}^{\mathsf{T}}\mathbf{X}$. Indeed, the eigenvector $\mathbf{z}$ associated with the eigenvalue $\gamma$ of the matrix $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ allows to write:
$\left(\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}\right)\mathbf{z}=\gamma\mathbf{z}\Rightarrow\mathbf{X}^{\mathsf{T}}\mathbf{X}\mathbf{z}=\left(\frac{\gamma-1}{\eta}\right)\mathbf{z}$
The above relation means the matrices $\mathbf{I}_{n}+\eta\mathbf{X}^{\mathsf{T}}\mathbf{X}$ and $\mathbf{X}^{\mathsf{T}}\mathbf{X}$ have the same eigenvectors which are associated with related eigenvalues, $j=1,2,\ldots,n$: